OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -negative one fourthx - 5. y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (anonymous):

@jigglypuff314 , @texaschic101

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Miss.SweetiePie

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

@campbell_st

OpenStudy (texaschic101):

first we find the slope of the equation given....y = -1/4x - 5 in y = mx + b form, the slope is in the m position y = (-1/4)x - 5 y = mx + b so what is the number in the m position....that is the slope

OpenStudy (texaschic101):

do you know the slope ?

OpenStudy (anonymous):

My guess was C, but I'm not sure.

OpenStudy (texaschic101):

we will get there....do you know the slope of the equation given ?

OpenStudy (anonymous):

No I am not sure.

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

ok....let me show you how to do this... the slope of the equation given is -1/4.....but we are looking for a perpendicular line...so we need the negative reciprocal slope. To find the negative reciprocal, flip the slope and change the sign. So the negative reciprocal of -1/4 is 4/1 (see how I flippped the slope and changed the sign) now we use y = mx + b our new slope (m) = 4 now we use the points given (-1,7)...x = -1 and y = 7) now we sub again y = mx + b 7 = 4(-1) + b 7 = -4 + b 7 + 4 = b 11 = b so our perpendicular equation is : y = 4x + 11.......your right, it is C

OpenStudy (anonymous):

Wow, thanks!

OpenStudy (texaschic101):

no problem :)